OpenStudy (anonymous):

When using the Cone Volume Formula can I turn 1/3 into 0.3? Formula ---> V = 1/3 x Pie(3.14) x R^2 x H

OpenStudy (paki):

1/3=0.333333333333333 so YES...

OpenStudy (anonymous):

yea u can

OpenStudy (paki):

@CrazyCountryGirl you are still here in the questions of cone volume :)

OpenStudy (anonymous):

Thank You! :0 I'll need you to check my answer in just a sec... :)

OpenStudy (paki):

sure

OpenStudy (anonymous):

Find the approximate volume of this cone. Use 3.14 for π. A. 37,680 cm^3 B. 56,520 cm^3 C. 113,040 cm^3 <--- D. 150,720 cm^3 Explanation... 0.3 x 3.14 x 900 x 40 / 3 = 113,040 @paki

OpenStudy (paki):

what is r and what is h...? values...?

OpenStudy (anonymous):

Oh sorry! Here you go.... :)

OpenStudy (paki):

3.14*900*40/3= 37,680 cm^3

OpenStudy (anonymous):

its A. 37,680 cm^3

OpenStudy (paki):

your above calculation is wrong......

OpenStudy (anonymous):

I see what I did wrong byt how come you didn't put in the 1/3 (or 0.3)?

OpenStudy (paki):

and why it is wrong.... you have putted down 1/3 and 0.3333 both values... use one....

OpenStudy (paki):

both values are the same... if you want to put 1/3 or 0.3333... both will give you the same result....

OpenStudy (paki):

@CrazyCountryGirl got it now...

OpenStudy (anonymous):

3.14*900*40/3= 37,680 cm^3 Where did you put the 1/3 in that equation? 0.3 x 3.14 x 900 x 40 / 3=113,040 I changed 1/3 to 0.3 and added it in the beginning. I thought the formula was V = 1/3 x 3.14 x R^2 x H

OpenStudy (paki):

yeah formula is V = 1/3 x 3.14 x R^2 x H... but you putted down 2 values of 1/3... have you noted that...?

OpenStudy (anonymous):

I get it! I divided in the 3 at the end and that's where the 1/3 goes?

OpenStudy (paki):

ok i am simplifying it more... see... you told me Where did you put the 1/3 in that equation... so see now... 3.14*900*40*1/3= 37,680 cm^3 ... clear now...

OpenStudy (anonymous):

Yes I get it now! Thanks so much! :)

OpenStudy (paki):

hmmm pleasure...

OpenStudy (paki):

@CO_oLBoY why you are silent bro.... :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

im not

OpenStudy (paki):

keep smiling yar.... :)

OpenStudy (anonymous):

i always do :)

OpenStudy (paki):

good

OpenStudy (anonymous):

yar koi achi movie to bta :P

OpenStudy (paki):

holiday 2014 :D

OpenStudy (anonymous):

ahan

OpenStudy (anonymous):

check krta

OpenStudy (paki):

han g check kar lo...

OpenStudy (anonymous):

yar english bta

OpenStudy (paki):

A Perfect Getaway (2009)...

OpenStudy (paki):

or 300.Rise.of.an.Empire.2014 or A Million Ways To Die In The West 2014

OpenStudy (paki):

aur douuuuuu :D

OpenStudy (anonymous):

de do :P

OpenStudy (anonymous):

mera konsa tax aa ra :P

OpenStudy (paki):

blended 2014

OpenStudy (paki):

aur agar BACHO wali daikhni hi tu RIO-2 download kar lo :D

OpenStudy (paki):

woh b 3D mai.... :D

OpenStudy (anonymous):

lol rehnw do :P

OpenStudy (paki):

theek hai

OpenStudy (anonymous):

btw anime me frozen or epic achi thin

OpenStudy (paki):

yar frozen achi hai...

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

despicable me bhi achi h XD